OpenStudy (deadshot):

How do I solve 64^(x) = 16^(x-1)?

OpenStudy (campbell_st):

rewrite them as powers of 4\[(4^3)^x = (4^2)^{x - 1}\] you can use the power of a power law to simplify them \[(x^a)^b = x^{a \times b}\] so once you have simplifed the powers, equate them and solve for x.

OpenStudy (deadshot):

So then it's x=-1, right?

OpenStudy (campbell_st):

nope...

OpenStudy (deadshot):

oh, i forgot to multiply the 2 and -1, so x=-2, right?

OpenStudy (campbell_st):

correct.

OpenStudy (deadshot):

Okay, thanks!